OpenStudy (anonymous):

Please help the equation of the curve is x lny = 2x + 1 the question asks to show that dy/dx = -y/x^2 i only got as far as : lny + x/y dy/dx = 2 now what ? thank you :D

geerky42 (geerky42):

Try divide both sides by x then take the derivative.

geerky42 (geerky42):

\[x \ln y = 2x + 1 \Rightarrow \ln y = \dfrac{2x + 1}{x}\] \[\dfrac{\text{d}}{\text{d}x}\ln y = \dfrac{\text{d}}{\text{d}x}\left(\dfrac{2x + 1}{x} \right) \] \[\dfrac{1}{y} \cdot \dfrac{\text{d}y}{\text{d}x} = \dfrac{\text{d}}{\text{d}x}\left(\dfrac{2x + 1}{x} \right) \]

OpenStudy (anonymous):

if i divide both sides by x i get 1/y dy/dx = 2/x - 2x -1

OpenStudy (anonymous):

ok so how do i get that to equal

OpenStudy (nubeer):

try looking at the steps @geerky42 have done.. your question is this. x lny = 2x + 1 just divide both sides by x

geerky42 (geerky42):

Maybe it won't work. All you have to do is to make equation clean and simple. From that, you probably will find a way to prove that dy/dx = -y/x².

geerky42 (geerky42):

Make it simple, then take the derivative of it then simplify it.

OpenStudy (nubeer):

hmm @geerky42 are u saying your method won't work the one u have mentioned or what?

geerky42 (geerky42):

I don't know. Try it and figure it out.

OpenStudy (nubeer):

it would work i have tried it..

geerky42 (geerky42):

If it doesn't work, find another way to approach to solution.

OpenStudy (anonymous):

nubeer you got -y/x^2 ?

OpenStudy (nubeer):

yup.. @geerky42 approach of solving the question is right.. just keep simplifying same way he has done.

geerky42 (geerky42):

Ok. Thanks. Do you know what to do now? @1234abdc

OpenStudy (anonymous):

no, I'm really confused

geerky42 (geerky42):

Just divide both sides by x so you won't have ugly and scary-look equation when you derivative it.

geerky42 (geerky42):

Do you know how to take the derivative, right?

OpenStudy (anonymous):

yes i do. my problem is simplifying

geerky42 (geerky42):

Well, can you do it from here?

OpenStudy (anonymous):

\[\frac{ 1 }{ y } \frac{ dy }{ dx } = \frac{ 2 }{ x } - 2x - 1\]

OpenStudy (anonymous):

now what ?

geerky42 (geerky42):

None, it's not correct. Do you know quotient rule, right?

OpenStudy (anonymous):

yes.. why is it wrong ?

geerky42 (geerky42):

I don't know how you got \( \dfrac{2}{x}-2x-1 \), but I got \( \dfrac{2x - (2x+1)}{x^2} \) which can be simplifed to \( -\dfrac{1}{x^2} \).

OpenStudy (anonymous):

oh ! i just got it i made a really stupid mistake thank you for your help :D

geerky42 (geerky42):

You're welcome. Glad I helped.

OpenStudy (help!!!!):

agreed